Tiago makes three models of camera lens. Its product mix and contribution margin per unit follow:

Percentage of Unit sales Contribution Margin per unit
Lens A 25 % $ 38
Lens B 40 30
Lens C 35 43

Required:

1. Determine the weighted-average contribution margin per unit.

2. Determine the number of units of each product that Tiago must sell to break even if fixed costs are $187,000.

3. Determine how many units of each product must be sold to generate a profit of $73,000.
Answer #1

Weighted average contribution margin

= (38*25%)+(30*40%)+(43*35%)

= 36.55

Breakeven point = Fixed cost/WA Contribution margin

= 187,000/36.55% = 5116

Units of A = 5116*25% = 1279

Units of B = 5116*40% = 2046

Units of C = 5116*35% = 1791

Breakeven point = (187,000+73,000)/36.55% = 7114

Units of A = 7114*25% = 1779

Uniys of B = 7114*40% = 2846

Units of C = 7114*35% = 2490